### v1

# Higher-Dimensional MOG dark compact object: shadow behaviour in the light of EHT observations

Abstract
Consideration of extra spatial dimensions is motivated by the unification of gravity with other interactions, the achievement of the ultimate framework of quantum gravity, and fundamental problems in particle physics and cosmology. Much attention has been focused on the effect of these extra dimensions on the modified theories of gravity. Analytically examining astrophysical phenomena like black hole shadows is one approach to understand how extra dimensions would affect the modified gravitational theories. The purpose of this study is to derive a higher-dimensional metric for a dark compact object in STVG theory and then examine the behaviour of the shadow shapes for this solution in STVG theory in higher dimensions. We apply the Carter method to formulate the geodesic equations and the Hamilton-Jacobi method to find photon orbits around this higher-dimensional MOG dark compact object. We investigate the effects of extra dimensions and the STVG parameter $\alpha$ on the black hole’s shadow size. Next, we compare the shadow radius of this higher-dimensional MOG dark compact object to the shadow size of the supermassive black hole M87*, which has been realized by the Event Horizon Telescope (EHT) collaborations, in order to restrict these parameters. We find that extra dimensions in the STVG theory typically lead to a reduction in the shadow size of the higher-dimensional MOG dark compact object, whereas the effect of parameter $\alpha$ on this black hole’s shadow is suppressible. Remarkably, given the constraints from EHT observations, we find that the shadow size of the four-dimensional MOG dark compact object lies in the confidence levels of the EHT data. Finally, we investigate the issue of acceleration bounds in higher-dimensional MOG dark compact object in confrontation with EHT data of M87*.
pacs:
04.50.Kd, 04.70.-s, 04.20.Ha, 04.25.dg, 04.50.-h, 04.50.Gh, 97.60.Lf
I Introduction
Dark compact objects, from a phenomenological perspective, represent a wide range of astronomical objects, such as white dwarfs, neutron stars, and black holes. From a theoretical perspective, it is possible to predict dark compact objects within the framework of extended gravity theories and in scenarios involving particle physics beyond the standard model Cardoso:2019rvt . Gravitational waves detected by LIGO/Virgo observations recently demonstrated the occurrence of binary black hole mergers, see LIGOScientific:2017bnn and references therein. In addition, the Event Horizon Telescope (EHT) demonstrated the presence of supermassive black holes in the Milky Way and M87 galactic centers EventHorizonTelescope:2021srq ; EventHorizonTelescope:2022xqj and references therein. Thus, it is not surprising that new species of compact objects will be discovered by future developments in the fields of very long baseline interferometry (VLBI) APhaseSensitiveInterferometer:1985 and gravitational wave astronomy. However, the study of dark compact objects from a mathematical perspective is interesting, because it can assist in comprehending how and at what limits a dark compact object tends to be a black hole by increasing its compactness. As mentioned, the Event Horizon Telescope (EHT) collaboration’s pioneering achievement of constructing the first pictures of the shadows of the supermassive black holes M87* (see EventHorizonTelescope:2021srq and references therein) and Sgr A* (see EventHorizonTelescope:2022xqj and references therein) has opened a new chapter in the physics of black holes and dark compact objects and enhanced our understanding of these mysterious celestial objects. Since a black hole does not emit light, its event horizon, that is, the surface at which matter or radiation cannot pass through, is not directly observable. Instead, what we can see is the ”shadow” of a black hole, a dark region surrounded by light that results in the gravitational lensing phenomenon around the event horizon  Perlick:2004tq ; Cosmpactness:2022 ; Relativisticimage:2009 ; ImagesDistortionHypothesis:2024 ; Cosmologicalconstant:2022 ; Ishihara:2016 ; Tsukamoto:2017 ; Ishihara:2017 ; Virbhadra:2000 ; Claudel:2001 ; Virbhadra:2002 ; Virbhadra:1998 ; Tsukamotoa:2023 ; Kumaran:2023 . Many efforts have been made to improve measurements in order to obtain high-resolution images since the release of the shadow images of M87* and Sgr A*  Goddi:2016qax . To achieve the needed resolution, theoretical efforts to study black hole physics, in particular, shadow behavior in different gravitational theories, has become of increasing significance. New theoretical shadow templates for upcoming observations are provided in this regard by analytical and numerical studies and analyses of the apparent geometrical shape of different black hole spacetimes. In addition to spacetime characteristics  Johannsen:2010ru ; Cunha:2017eoe and the observer’s position, the black hole parameters; mass, electric charge, and angular momentum determine the shadow’s shape and size deVries:1999tiy . The shadow shape for non-rotating black holes is a perfect circle. Rotating black holes can have non-trivial shadow patterns due to the angular momentum parameter  Chandrasekhar:1984siy .
Even though Albert Einstein’s General Theory of Relativity (GR) has made significant progress in explaining observations and predicting astounding events, it is still unable to fully explain gravitational interaction and the events that correspond with it in the universe. Some instances of the shortcomings of GR are the reproduction of the rotation curves of nearby galaxies Sofue:2000jx ; Sofue2016 , mass profiles of galaxy clusters Ettori2013 ; Voigt2006 , inherent singularities at the center of black holes, etc. Furthermore, the positively accelerated expansion of the Universe at late time is explained by GR only if the cosmological constant factor $\Lambda$ is present Weinberg1989 ; Peebles2003 . Restructuring the geometric part of General Relativity (GR) using various ways is an intriguing reform strategy. One such approach is the so-called MOdified Gravity (MOG), a Scalar-Tensor-Vector (STVG) theory developed and proposed by John W. Moffat to describe gravitational interaction Moffat2006 . The gravitational effects of spacetime in the MOG setup are expressed by a massive vector field $\phi_{\mu}$ in addition to three scalar fields: the mass of the vector field $\tilde{\mu}$, the effective gravitational constant $G$, and the vector field coupling $\xi$. In addition to being compatible with cosmological observational data Davari2021 and Planck 2018 data  Moffat2021a , the MOG theory has achieved various successes in explaining astrophysical phenomena. These include explaining the dynamics of galaxy clusters without dark matter and the rotation curves of many galaxies  Brownstein2006a ; Brownstein2006b ; Brownstein2007 ; Moffat2013 ; Moffat2014 ; Moffat2015a . Furthermore, the framework of MOG theory is currently experiencing the release of a number of black hole solutions, including non-rotating and rotating ones  Moffat2015b , even with extra dimensions Cai2021 , cosmological solutions Roshan2015 ; Jamali2018 ; Davari2021 , and non-stationary solutions for inhomogeneity distributions of mass–energy in spacetime  Perez2019 . Additionally, a great deal of theoretical and observational work has been done to comprehend the characteristics of the MOG theory and how it functions in various contexts  Moffat2021b ; Guo2018 ; Moffat2020 ; Mureika2016 ; Saghafi2021 ; Saghafi:2021wzx ; Perez:2017spz ; Shukla:2022sti ; Hussain:2015cga ; Sharif:2017owq ; Lee:2017fbq ; Kolos:2020ykz ; Hu:2022lek ; John:2019was . Ref.  Moffat:2018jmi offers an interesting exploration of the solution characterizing the regular rotating and non-rotating MOG dark compact object. Ref.  Sau:2022afl examines the shadow behavior of the regular rotating and non-rotating MOG dark compact object.
The notion of extra dimensions was initially put forward by Nordstrom in 1914  Nordstrom:1914ejq . His theory states that by treating four-dimensional spacetime as a surface in a five-dimensional spacetime, one may unify the gravitational and electromagnetic fields. Currently, the primary motivations of the vast amount of research on the extra dimensions are the quantization of gravitational interaction, the unification of gauge and gravitational interactions of elementary particles, the Higgs mass hierarchy problem, and the cosmological constant problem. This is why the Kaluza-Klein (KK) theory Kaluza:1921tu ; Klein:1926tv unifies electromagnetic or even non-Abelian gauge fields that characterize weak and strong interactions with gravitational interaction by introducing a compact space constructed by compact extra dimensions with a specific compactification scale. Additionally, as a well-known candidate theory for quantum gravity, the string theory (M-theory) has ten or eleven compact dimensions Witten:1995ex ; Schwarz:1995jq . There are ideas for large additional dimensions of the order of millimeter in addition to these compact extra dimensions with the extension up to the order of the Planck length. The Arkani-Hamed-Dimopoulos-Dvali (ADD) braneworld model Arkani-Hamed:1998jmv ; Arkani-Hamed:1998sfv has provided a novel path for the study of extra dimensions by using large extra dimensions to solve the Higgs mass hierarchy problem. It is important to remember that the striking aspect of these large extra dimensions is that future accelerator, astrophysical, and even tabletop studies will be able to measure their effects. It is surprising to learn that string theory can incorporate the ADD model  Antoniadis:1998ig . The Randall-Sundrum (RS) braneworld model Randall:1999ee ; Randall:1999vf proposes the use of warped extra dimensions in addition to compact and large extra dimensions to solve the Higgs mass hierarchy problem. Apart from the aforementioned extra dimensions, there exist theories with unlimited volume extra dimensions such as the Dvali-Gabadadze-Porrati (DGP) braneworld scenario  Dvali:2000hr , wherein the spacetime is not four-dimensional even at very low energies and the extra dimensions are neither wrapped nor compact. These theories, which alter gravity at large distances because the existence of infinite volume additional dimensions, are possibilities for solving the cosmological constant problem Dvali:2000xg ; Dvali:2002pe . Refs. Gabadadze:2003ii ; Shifman:2009df ; Perez-Lorenzana:2005fzz ; Rubakov:2001kp provide some thorough evaluations of higher-dimensional models. A variety of techniques and strategies have been used within the context of black hole physics to expand and study alternative black hole models in arbitrary dimensions Emparan:2008eg ; Kanti:2004nr . One such technique is the Tangherlini method Tangherlini:1963bw , which generalizes the Schwarzschild solution to $n$ dimensions.
In high-energy or particle experiments, physicists concentrate on detecting extra dimensions. Prospective instruments for investigating the impact of strong gravity regimes corresponding to higher-dimensional black holes and extra dimensions are the Large Hadron Collider (LHC) at CERN and future colliders Allanach:2002gn ; Agashe:2006hk ; Franceschini:2011wr ; Deutschmann:2017bth ; Strominger:1996sh ; Harris:2004xt ; Antoniadis:2000vd . Furthermore, evidence for the existence of extra dimensions comes from spectroscopic tests Zhou:2014xbw ; Luo:2006ck ; Luo:2006ad , the presence of hydrogen atoms in higher dimensions Burgbacher:1999sha ; Caruso:2012daf ; Shaqqor:2009cha , and theories to solve the proton radius puzzle Wang:2013fma ; Dahia:2015bza ; Zhi-gang:2007swh . However, the detection of gravitational waves (GW) by the LIGO/Virgo collaborations LIGOScientific:2016aoc and the aforementioned images obtained by the EHT collaborations are two recent successes in the search for black hole strong field regimes. Some traces of additional dimensions exist in the GW detection, with specific details regarding the associated amplitude and fluctuation mode dynamics. As a result, a lot of research has been done to uncover this physics Cardoso:2016rao ; Yu:2016tar ; Visinelli:2017bny ; Kwon:2019gsa ; for a thorough analysis, refer to Ref. Yu:2019jlb . EHT has now opened up novel opportunities for the search for extra dimensions. The authors of recent seminal publications Vagnozzi:2019apd ; Belhaj:2020mlv ; Tang:2022hsu discovered significant limitations on warped and compact extra dimensions within the RS model and M-theory, derived from EHT observations. Thus, as we aim to do in this study, one can use the EHT data to investigate all kinds of extra dimensions in general and see if they can be detected. In this sense, it appears that extra dimensions have an impact on black hole shadows by decreasing the size of the shadow in a variety of black hole models and gravitational theories Amarilla:2011fx ; Eiroa:2017uuq ; Papnoi:2014aaa ; Singh:2017vfr ; Amir:2017slq ; Belhaj:2020rdb . Nevertheless, there is still much to learn about the precise impact of extra dimensions on black hole shadows, which is the subject of ongoing research. It appears that there is a greater chance of future detection of the effects of large and infinite volume extra dimensions.
Furthermore, due to extra degrees of freedom that arose from extended theories of gravity, black hole shadow size and shape may vary. Thus, determining the size and shape of black hole shadows could help evaluate black hole metrics parameters and test alternative theories of gravity. The question of black hole shadow has been the subject of a great deal of research to determine whether and how degrees of freedom—derived from extended theories of gravity other than black hole parameters—affect the behavior of the shadow Perlick:2021aok . Here are a few instances: Black hole shadows in Einstein-Maxwell-dilaton gravity Amarilla:2013sj ; Wei:2013kza , in Chern-Simons modified gravity Amarilla:2010zq , the shadow behavior of the Kerr-Newman family of solutions of the Einstein-Maxwell equations is studied in Refs. Bardeen:1973tla ; Takahashi:2004xh ; the shadow of a black hole with NUT-charges Chakraborty:2013kza ; Grenzebach:2014fha ; the apparent shape of the Sen black hole  Hioki:2008zw ; Dastan:2016bfy ; Younsi:2016azx ; shadow behavior of the non-commutative geometry inspired and quantum-corrected black holes  Wei:2015dua ; Sharif:2016znpx ; Saghafi:2022pmey ; shadow behavior of Einstein-Born-Infeld black holes  Atamurotov:2015xfa ; shadow behavior of Ayon-Beato-Garcia black hole and also, rotating Hayward and rotating Bardeen regular black holes  Abdujabbarov:2016hnw and hairy black holes Cunha:2015yba ; Cunha:2016bjh ; shadows of colliding and multi-black holes Nitta:2011nin ; Yumoto:2012kz ; shadow behavior of rotating black holes in $f(R)$ gravity Dastan:2016vhb , conformal Weyl gravity  Mureika:2016efo , and Einstein-dilaton-Gauss-Bonnet black holes  Cunha:2016wzk ; Shadow behavior of wormholes and naked singularities Nedkova:2013msa ; Ohgami:2015nra ; Ortiz:2015rma , as well as chaotic shadow of a non-Kerr rotating compact object with quadrupole mass moment and a magnetic dipole  Wang:2018eui ; Wang:2017qhh , and black holes with exotic matter  Tinchev:2015apf ; Abdujabbarov:2015pqp ; Singh:2017xle ; Huang:2016qnl .
In this article, we aim to develop a higher-dimensional metric for a dark compact object in STVG theory, derive its black hole solutions, and investigate its horizons in different dimensions. Afterward, we explore the optical properties of dark compact object, such as deflection, shadow size, photon sphere, and effective potential ($V_{eff}$). Finally by calculating the deflection angle and shadow size in different dimensions, and comparing them with observational data obtained from M87* and Sgr A*, we evaluate whether our universe is an extra dimensional universe. Additionally, we will assess the level of acceptance of STVG theory as an alternative gravitational theory.
This paper is organized as follows. In section II we first introduce the Modified Gravity theory and then introduce the line element of the higher-dimensional dark compact object in this theory. In section III we study the shadow behavior, energy emission rate, and deflection angle of the STVG dark compact object in MOG theory with extra dimensions, and we analyze the impact of parameter $\alpha$ and dimensions on the shadow and deflection angle of the dark compact object. In section IV we constrain the parameter $\alpha$ by comparing the shadow size of the suppermassive black hole M87* with the shadow size of the STVG dark compact object. In section V we examin the issue of acceleration bounds in higher-dimensional MOG dark compact object’s spacetime background. Finally, Section VI is devoted to summary, conclusion and discussing about our main results.
II MOG and its higher-dimensional regular static spherically symmetric dark compact object
The STVG theory’s overall action takes the form of Moffat2006 :
$S=S_{GR}+S_{M}+S_{\phi}+S_{S}\,,$ (1)
where $S_{GR}$ is the Einstein-Hilbert action, $S_{M}$ is the action of all possible matter sources, $S_{\phi}$ is the action of the (spin $1$ graviton) vector field $\phi^{\mu}$ possessing the mass $\tilde{\mu}$ as one of the scalar fields in the theory, and $S_{S}$ is the action of three scalar fields, which can be expressed as follows
$S_{GR}=\frac{1}{16\pi}\int d^{D}x\sqrt{-g}\frac{1}{G}R\,,$ (2)
$S_{\phi}=-\int d^{D}x\sqrt{-g}\left(\frac{1}{4}B^{\mu\nu}B_{\mu\nu}+V_{1}(\phi)\right)\xi\,,$ (3)
$\begin{split}S_{S}&=\int d^{D}x\sqrt{-g}\left[\frac{1}{G^{3}}\left(\frac{1}{2}g^{\mu\nu}\nabla_{\mu}G\nabla_{\nu}G-V_{2}(G)\right)+\frac{1}{\tilde{\mu}^{2}G}\left(\frac{1}{2}g^{\mu\nu}\nabla_{\mu}\tilde{\mu}\nabla_{\nu}\tilde{\mu}-V_{3}(\tilde{\mu})\right)\right.\\ &\left.+\frac{1}{G}\left(\frac{1}{2}g^{\mu\nu}\nabla_{\mu}\xi\nabla_{\nu}\xi-V_{4}(\xi)\right)\right]\,,\end{split}$ (4)
in which $D$ represents the dimension of spacetime, $g_{\mu\nu}$ is the background metric tensor and $g$ is the corresponding determinant, $R$ is the Ricci scalar constructed by contracting $R_{\mu\nu}$ as the Ricci tensor, $G$ is a scalar field in the setup, which known as the enhanced Newtonian parameter, $\xi$ is third scalar field in the setup as the vector field coupling, $V_{1}(\phi)$, $V_{2}(G)$, $V_{3}(\tilde{\mu})$, and $V_{4}(\xi)$ are the corresponding potentials of the vector field $\phi^{\mu}$, and three scalar fields $G$, $\tilde{\mu}$, and $\xi$, respectively, and $B_{\mu\nu}=\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}$, and also $\nabla_{\mu}$ stands for the covariant derivative in the spacetime.
In the STVG theory, $T_{\mu\nu}={}^{(M)}T_{\mu\nu}+{}^{(\phi)}T_{\mu\nu}+{}^{(S)}T_{\mu\nu}$ is the total stress-energy tensor, in which the stress-energy tensor of matter sources is ${}^{(M)}T_{\mu\nu}$, the stress-energy tensor of the scalar fields is ${}^{(S)}T_{\mu\nu}$, and the stress-energy tensor of the vector field is
${}^{(\phi)}T_{\mu\nu}=-\frac{1}{4}\left(B_{\mu}^{\,\,\,\sigma}B_{\nu\sigma}-\frac{1}{4}g_{\mu\nu}B^{\sigma\lambda}B_{\sigma\lambda}\right)\,,$ (5)
for which $V_{1}(\phi)=0$. One can find the full field equations of the STVG framework by variation of the action $S$ concerning the inverse of the metric tensor, which yields Moffat2006
$G_{\mu\nu}+G\left(\nabla^{\gamma}\nabla_{\gamma}\frac{1}{G}g_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}\frac{1}{G}\right)=8\pi GT_{\mu\nu}\,,$ (6)
in which the Einstein tensor is defied as $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$ and we have set $c=1$.
II.1 Regular MOG static spherically symmetric dark compact object
The line element of the regular MOG static spherically symmetric dark compact object were found under the following assumptions Moffat:2018jmi
  • The vector field is massless, i.e., $\tilde{\mu}=0$, since one can prove that for MOG compact objects, e.g., black holes possessing horizons, the mass of the vector field in the setup is zero.
  • The enhanced Newtonian parameter $G$ is defined as a constant depending on the free dimensionless parameter $\alpha$ so that $G=G_{N}(1+\alpha)$ where $G_{N}$ is the Newtonian constant. Furthermore, the gravitational source charge of the vector field is $Q_{g}=\sqrt{\alpha G_{N}}M$ where $M$ is the source mass. Here, we set $G_{N}=1$.
  • The vector field coupling is set to unity, i.e., $\xi=1$.
  • The matter-free field equations of STVG setup is considered since the MOG dark compact object is a vacuum solution of the framework.
The above assumptions result in $S_{M}=S_{S}=0$ and consequently, we have ${}^{(M)}T_{\mu\nu}={}^{(S)}T_{\mu\nu}=0$. Thus, the field equations (6) now reduce to the following form
$G_{\mu\nu}=8\pi(1+\alpha){}^{(\phi)}T_{\mu\nu}\,.$ (7)
To solve a higher-dimensional static spherically symmetric compact object in the STVG, the line element is typically assumed to have the following form:
$ds^{2}=-h(r)dt^{2}+\frac{1}{f(r)}dr^{2}-r^{2}d\Omega_{D-2}^{2}\,,$ (8)
where $d\Omega^{2}_{D-2}=d\theta_{1}^{2}+\sin^{2}\theta_{1}d\theta_{1}^{2}+...+\prod_{i=1}^{D-3}\sin^{2}\theta_{i}d\theta_{D-2}^{2}$ represents the line element on the $(D-2)$-dimensional unit sphere.
$f(r)=h(r)=1-\frac{(D-2)mr^{2(D-3)}\omega_{D-2}}{8\pi(r^{2(D-3)}+\frac{(D-2)^{2}m^{2}\alpha(1+\alpha)\omega^{2}_{D-2}}{256G^{2}\pi^{2}})^{\frac{3}{2}}}+\frac{(D-3)(D-2)Gq^{2}r^{2(D-3)}\omega^{2}_{D-2}}{32\pi^{2}(r^{2(D-3)}+\frac{(D-2)^{2}m^{2}\alpha(1+\alpha)\omega^{2}_{D-2}}{256G^{2}\pi^{2}})^{2}}$ (9)
where $m$ and $q$ are defined by
$m\equiv\frac{16\pi GM}{(D-2)\omega_{{}_{D-2}}},\qquad q\equiv\frac{8\pi Q}{\sqrt{2(D-2)(D-3)}\omega_{{}_{D-2}}},$ (10)
and $\omega_{{}_{D-2}}=\frac{2\pi^{\frac{D-1}{2}}}{\Gamma(\frac{D-1}{2})}$ means the area of a $(D-2)$-dimensional unit sphere.
It is obvious that the metric function Eq. (9) turns back to the MOG dark compact object solution given by Moffat Moffat:2018jmi for the case of $D=4$. The MOG dark compact object possesses a critical value for $\alpha$ as $\alpha_{crit}=0.674$ Moffat:2018jmi , so that for $\alpha\leq\alpha_{crit}$ it has two horizons. It is worth mentioning that the (spin $1$ graviton) vector field produces a repulsive gravitational force, which prevents the collapse of the MOG dark compact object to a MOG black hole with horizon.
Setting $\alpha=0$ in the line element (8) recovers the Schwarzschild-Tangherlini black hole in GR. Moreover, the asymptotic behavior of the higher-dimensional MOG compact object in the limit of $r\rightarrow\infty$ is deduced as follows
$f(r)\approx 1-\frac{2(1+\alpha)m}{r^{D-3}}+\frac{\alpha(1+\alpha)Gq^{2}}{r^{2(D-3)}}\,.$ (11)
When, $\alpha\leq\alpha_{crit}$, the two horizons of the regular higher-dimensional MOG static spherically symmetric dark compact object in the limit of $r\rightarrow\infty$ can be found as
$r_{\pm}=\Big{(}M+M\alpha\pm\sqrt{M^{2}+M^{2}\alpha}\Big{)}^{\frac{1}{D-3}}\,.$ (12)
Where $r_{-}$ is the inner horizon called the Cauchy horizon and $r_{+}$ is the outer horizon called the event horizon. In addition, when $\alpha=0$, the two horizons are merged into the event horizon of the Schwarzschild-Tangherlini black hole Tangherlini:1963 ; polchinski:1998 . When, $\alpha>\alpha_{crit}$, there is a naked regular MOG static spherically symmetric dark compact object with no horizon. On the other hand, approaching the source, i.e., $r\rightarrow 0$, the 4D-MOG dark compact object behaves to the form
$f(r)\approx 1-\frac{r^{2}}{M^{2}}\left(\frac{2\sqrt{1+\alpha}-\sqrt{\alpha}}{(1+\alpha)\alpha^{\frac{3}{2}}}\right)\,.$ (13)
Therefore, the spacetime metric of the MOG dark compact object is regular so that $f(0)=1$. Additionally, one can verify that the Kretschmann scalar $R^{\mu\nu\lambda\sigma}R_{\mu\nu\lambda\sigma}$ in addition to the Ricci scalar $R$ in the spacetime metric are regular at $r=0$.
For the static spherically symmetric system, the gravitational redshift $z$ at the asymptotic distance $r$ to an observer is gathered as follows
$z(r)=\frac{1}{\sqrt{f(R)}}-1\,,$ (14)
where the MOG dark compact object’s radius is $R$. The gravitational redshift of the compact object becomes infinite on the horizon $r_{+}$ for $\alpha<\alpha_{crit}$, and has a finite value for $\alpha>\alpha_{crit}$ in the limit of $r\rightarrow\infty$. However, one expects the regular MOG dark compact object in order to be sufficiently dark in order to be consistent with binary X-ray observations based on observational data., so that $\alpha\sim\alpha_{crit}$ Moffat:2018jmi .
In spacetime dimensions $D=4,5,6,7$, we plot the graph of the metric function $f(r)$ with respect to $r$ for various values of $\alpha$. It should be noted that the Schwarzschild-Tangherlini black hole case in Einstein’s gravity is represented by the parameter $\alpha=0$. This image illustrates how, for $D=4,5,6,7$, the event horizon radius $r_{\rm H}$ decreases as the parameter $\alpha$ decreases. Furthermore, before we reach the extreme situation, we may observe that the higher-dimensional MOG dark compact object always has one more horizon than the Schwarzschild-Tangherlini black hole. Furthermore, by increasing the spacetime dimension both the event horizon and the inner Cauchy horizon radius decrease.
Figure 1: Function $f(r)$ with respect to $r$ for different values of $\alpha$, where the spacetime dimensions are taken to be $D=4,5,6,7$, respectively. Here we set $M=1$.
III Shadow behaviour of the higher-dimensional MOG dark compact object
A part of the light that would reach the observer is deflected by the black hole’s gravitational field when it passes in front of it. The apparent shape of the black hole is the boundary of the shadow, which is created when photons fall into the black hole and form a dark area known as the shadow. The general formulae needed to determine the deflection angle, energy emission rate, and shadow shape for the higher-dimensional general ansatz (8)—which requires examining a test particle’s motion in spacetime—are provided in this section. Then we apply higher-dimensional MOG dark compact object line element on the formulas deduced in this section.
In MOG theory with extra dimensions, we analyze the shadow and deflection angle of the dark compact object by taking arbitrary values for the STVG parameter, $\alpha=0,0.2,0.45,0.5$, and $0.674$. Furthermore, we consider that the additional dimensions count $D=4,5,...,11$ (keep in mind that $D=4$ represent one temporal in addition to three spatial dimensions as usaul). In Sau:2022afl , the authors consider regular MOG dark compact objects, using nonlinear field equations to determine the effects of $\phi_{\mu}$ for $\alpha>0.674$ by adopting the ADM mass, here we examine the case for $\alpha\leq\alpha_{crit}$.
III.1 Effective potential
Initially, we aim to examine the effective potential behaviour of the dark compact object with arbitrary dimensions. To do this, we start with the Lagrangian of the test particle that has the following form
~$\tilde{\mathcal{L}}=\frac{1}{2}g_{ab}\dot{x}^{a}\dot{x}^{b}\,,$ (15)
in which dot is the derivative with respect to affine parameter $\tau$. The canonically conjugate momentum components corresponding with the Eq. (8) are as follows
$P_{t}=h(r)\dot{t}=E\,,$ (16)
$P_{r}=\frac{1}{f(r)}\dot{r}\,,$ (17)
$P_{\theta_{i}}=r^{2}\sum_{i=1}^{D-3}\prod_{D=1}^{i-1}\sin^{2}(\theta_{D})\dot{\theta}_{i}\,,$ (18)
$P_{\theta_{D-2}}=r^{2}\prod_{i=1}^{D-3}\sin^{2}(\theta_{i})\dot{\theta}_{D-2}=L\,,$ (19)
where $i=1,2,3,...,D-3$, and $L$ is the angular momentum and $E$ is the energy of the test particle. We apply the Carter approach to study the geodesic equations and the Hamilton-Jacobi method to analyze photon orbits around the black hole. We need to generalize these methods to higher dimensions for our work. The Hamilton-Jacobi method in extra dimensions has the following form
$\frac{\partial S}{\partial\tau}=-\frac{1}{2}g^{ab}\frac{\partial S}{\partial x^{a}}\frac{\partial S}{\partial x^{b}}\,,$ (20)
in which $S$ is the Jacobi action of the test particle. Inserting the general ansatz. (8) with arbitrary dimensions into Eq. (20), we have
$\begin{split}-2\frac{\partial S}{\partial\tau}=-\frac{1}{h(r)}\left(\frac{\partial S_{t}}{\partial t}\right)^{2}+f(r)\left(\frac{\partial S_{r}}{\partial r}\right)^{2}+\sum_{i=1}^{D-3}\frac{1}{\left(r^{2}\prod_{D=1}^{i-1}\sin^{2}\theta_{D}\right)}\left(\frac{\partial S_{\theta_{i}}}{\partial\theta_{i}}\right)^{2}+\frac{1}{\left(r^{2}\prod_{i=1}^{D-3}\sin^{2}\theta_{i}\right)}\left(\frac{\partial S_{\theta_{D-2}}}{\partial\theta_{D-2}}\right)^{2}\,.\end{split}$ (21)
By considering a separable solution for Jacobi action, the action can be written as
$S=\frac{1}{2}m^{2}\tau-Et+L\theta_{D-2}+S_{r}(r)+\sum^{D-3}_{i=1}S_{\theta_{i}}(\theta_{i})\,,$ (22)
where $m$ is the rest mass of the test particle. Since the photon is the test particle used to analyze the shadow behavior of black holes, we set $m=0$. Also by inserting the action (22) into Eq. (21) we derive the following equations
$\begin{split}0&=\left\{\frac{E^{2}}{h(r)}-f(r)\left(\frac{\partial S_{r}}{\partial r}\right)^{2}-\frac{1}{r^{2}}\left(\frac{L^{2}}{\prod_{i=1}^{D-3}\sin^{2}\theta_{i}}+\mathcal{K}-\prod_{i=1}^{D-3}L^{2}\cot^{2}\theta_{i}\right)\right\}\\ &-\left\{\frac{1}{r^{2}}\left(\sum_{i=1}^{D-3}\frac{1}{\prod_{D=1}^{i-1}\sin^{2}\theta_{D}}\left(\frac{\partial S_{\theta_{i}}}{\partial\theta_{i}}\right)^{2}-\mathcal{K}+\prod_{i=1}^{D-3}L^{2}\cot^{2}\theta_{i}\right)\right\}\,,\end{split}$ (23)
where $\mathcal{K}$ is the Carter constant. The following set of equations can be obtained after some manipulations
$r^{4}f^{2}(r)\left(\frac{\partial S_{r}}{\partial r}\right)^{2}=r^{4}\frac{f(r)}{h(r)}E^{2}-r^{2}\left(L^{2}+\mathcal{K}\right)f(r)\,,$ (24)
$\sum_{i=1}^{D-3}\frac{1}{\prod_{D=1}^{i-1}\sin^{2}\theta_{D}}\left(\frac{\partial S_{\theta_{i}}}{\partial\theta_{i}}\right)^{2}=\mathcal{K}-\prod_{i=1}^{D-3}L^{2}\cot^{2}\theta_{i}\,.$ (25)
Finally, the equation of motion for photon (null geodesic within the higher-dimentional spacetime (8)) can be derived as follows, by utilizing Eqs. (24) and the canonically conjugate momentum components (16)-(19)
$\dot{t}=\frac{E}{f(r)}\,,$ (26)
$r^{2}\dot{r}=\pm\sqrt{\mathcal{R}}\,,$ (27)
$r^{2}\sum_{i=1}^{D-3}\prod_{D=1}^{i-1}\sin^{2}(\theta_{D})\dot{\theta}_{i}=\pm\sqrt{\Theta_{i}}\,,$ (28)
$\dot{\theta}_{D-2}=\frac{L}{r^{2}\prod_{i=1}^{D-3}\sin^{2}\theta_{i}}\,,$ (29)
where “$+$” and “--” signs denote the outgoing and ingoing radial directions of the motion of photon, respectively. Also we have the motion of photon in the spacetime as follows which can be derived by Eqs. (26)-(29).
$\mathcal{R}=r^{4}\frac{f(r)}{h(r)}E^{2}-r^{2}\left(L^{2}+\mathcal{K}\right)f(r)\,,\quad\Theta_{i}=\mathcal{K}-\prod_{i=1}^{D-3}L^{2}\cot^{2}\theta_{i}\,.$ (30)
Rewriting the radial null geodesic equation (27) yields the effective potential,which has the following form:
$V_{eff}=\frac{f(r)}{r^{2}}(\mathcal{K}+L^{2})-\frac{f(r)}{h(r)}E^{2}$ (31)
where $\mathcal{K}$, $E$ and $L$ are the constants of motion.
The unstable circular orbits of photons define the boundary of apparent shape of the black hole. They correspond with the maximum value of the effective potential, which occurs at a distance, referred as photon sphere radius $r_{0}$ satisfyig the equations below
$V_{eff}\big{|}_{r_{0}}=\frac{dV_{eff}}{dr}\bigg{|}_{r_{0}}=0\,,\quad\mathcal{R}\big{|}_{r_{0}}=\frac{d\mathcal{R}}{dr}\bigg{|}_{r_{0}}=0\,.$ (32)
Therefore, the least value of the roots of the following equation is the photon sphere radius $r_{0}$ associated with the maximum of the black hole’s effective potential in spacetime (8) with arbitrary dimensions,
$r_{0}h^{\prime}(r_{0})-2h(r_{0})=0\,,$ (33)
where a prime stands for radial derivative.
By inserting Eq. (9) into Eq. (31) we can derive the effective potential of higher-dimensional MOG dark compact object as follows
$V_{eff}=1-\frac{(D-2)mr^{2(D-4)}\omega_{{}_{D-2}}}{8\pi(r^{2(D-3)}+\frac{(D-2)^{2}m^{2}\alpha(1+\alpha)\omega^{2}_{D-2}}{256G^{2}\pi^{2}})^{\frac{3}{2}}}+\frac{(D-3)(D-2)Gq^{2}r^{2(D-4)}\omega^{2}_{D-2}}{32\pi^{2}(r^{2(D-3)}+\frac{(D-2)^{2}m^{2}\alpha(1+\alpha)\omega^{2}_{D-2}}{256G^{2}\pi^{2}})^{2}}(\mathcal{K}+L^{2})-E^{2}$ (34)
(a) for $\alpha=0.5$
(b) for $D=4$
(c) for $D=6$
Figure 2: The figure of the radial variation of effective potential of the dark compact object with extra dimensions in STVG theory for various values of $D$ and $\alpha$,where we set $L=5$ and $E=\mathcal{K}=M=1$.
For various values of $D$ and $\alpha$, the behavior of the effective potential for the higher dimensional MOG dark compact object is shown in Fig. 2 as a function of radial coordinate $r$. The photon sphere radius $r_{0}$, which corresponds to each value of $D$ and $\alpha$, is where the effective potential in this figure peaks, and as $r\rightarrow\infty$, the effective potential approaches a steady value. As we can observe from Fig. 2a, the higher-dimensional MOG dark compact object’s effective potential grows as $D$ increases for a constant value of $\alpha$. Furthermore, we discover from Figs. 2b and 2c that raising $\alpha$ for $D=4,6$ causes the dark compact object’s effective potential to be amplified. By examining these figures it is obvious that the impact of higher dimensions dominates the effect of $\alpha$ in the MOG dark compact object by increasing the spacetime dimensionality. Figure 2 illustrates how $D$ and $\alpha$ impact the shadow boundary of the MOG dark compact object by increasing the spacetime dimensionality. This is because the location of the black hole’s maximum effective potential, or the photon sphere radius $r_{0}$, describes the black hole’s shadow boundary.
III.2 Geometrical shapes of shadow
We want to determine the shadow form and size of the black hole in the arbitrary-dimensional spacetime (8) in this section. First, we define two impact parameters, $\xi$ and $\eta$, in order to accomplish this end. The characteristics of photons close to black holes can be described by these impact parameters as functions of the constants of motion $E$, $L$, and $\mathcal{K}$. They provide the following definitions:
$\xi=\frac{L}{E}\,,\qquad\eta=\frac{\mathcal{K}}{E^{2}}\,.$ (35)
Therefore, one can rewrite the effective potential and also, the function $\mathcal{R}$ in terms of these impact parameters as
$V_{eff}=E^{2}\left\{\frac{f(r)}{r^{2}}\left(\eta+\xi^{2}\right)-\frac{f(r)}{h(r)}\right\}\,,\qquad\mathcal{R}=E^{2}\left\{r^{4}\frac{f(r)}{h(r)}-r^{2}f(r)\left(\eta+\xi^{2}\right)\right\}\,.$ (36)
Finally, by inserting Eq. (36) into Eq. (32) one can find the following equation for two unknowns $\xi$ and $\eta$
$\eta+\xi^{2}=\frac{r_{0}^{2}}{2f(r_{0})+r_{0}f^{\prime}(r_{0})}\left\{4\left(\frac{f(r_{0})}{h(r_{0})}\right)+r_{0}\left(\frac{f^{\prime}(r_{0})h(r_{0})-f(r_{0})h^{\prime}(r_{0})}{h^{2}(r_{0})}\right)\right\}\,.$ (37)
Consequently, using Eq. (37), the photon sphere radius obtained from Eq. (33) yields the quantity $\eta+\xi^{2}$. It is evident that the quantity $\eta+\xi^{2}$ has the dimension of the length square, and $r_{0}$ has the dimension of the length.
Usually $\lambda$ and $\psi$ are used as celestial coordinates to describe the geometric shape of the shadow as perceived by the observer’s frame Vazquez:2003zm . These coordinates can be read as follows
$\lambda=\lim_{r_{o}\rightarrow\infty}\left(\frac{r_{o}^{2}P^{(\theta_{n-2})}}{P^{(t)}}\right)\,,\qquad\psi=\lim_{r_{o}\rightarrow\infty}\left(\frac{r_{o}^{2}P^{(\theta_{i})}}{P^{(t)}}\right)\,,$ (38)
where $\left[P^{(t)},P^{(\theta_{D-2})},P^{(\theta_{i})}\right]$ are its vi-tetrad momentum elements and $r_{o}$ is the distance between the observer and the black hole. $\psi=\pm\sqrt{\eta}$ and $\lambda=-\xi$ are found on the equatorial plane. As a result, we can achieve the following result
$R_{s}^{2}\equiv\eta+\xi^{2}=\lambda^{2}+\psi^{2}\,,$ (39)
where the shadow radius in celestial coordinates is denoted by $R_{s}$. The geometric shape of the shadow for non-rotating (static) black holes is a circle of radius $R_{s}$.
We will now demonstrate the geometrical shape of the MOG dark compact object’s shadow in extra dimensions on the observer’s sky using the celestial coordinates that were presented above. To begin with, we should collect some numerical data for the black hole’s related $r_{eh},r_{0}$, and $\sqrt{\eta+\xi^{2}}$. For the higher-dimensional MOG dark compact object, the photon sphere radius may be obtained by inserting Eq. (9) into Eq. (33). Furthermore, the radius of shadow circles for the higher-dimensional MOG dark copact object in celestial coordinates may be found by utilizing Eq. (39) and applying Eq. (9) into Eq. (37). In Table 1 we collect the numerical data associated with $r_{e}h$, and $r_{0}$ for $D=4,5,6,7,8$ and some different values of $\alpha$.
Table 1: Values of $r_{eh}$, and $r_{0}$ for different values of $\alpha$ and $D$.
$D=4$ $D=5$ $D=6$ $D=7$ $D=8$
$D$ $r_{eh}$ $r_{0}$ $r_{eh}$ $r_{0}$ $r_{eh}$ $r_{0}$ $r_{eh}$ $r_{0}$ $r_{eh}$ $r_{0}$
$\alpha=0$ 2 3 1.4142 2 1.25992 1.70998 1.1892 1.56508 1.14869 1.58482
$\alpha=0.2$ 2.29545 3.26804 1.51507 2.11584 1.31913 1.78537 1.23088 1.62095 1.18079 1.61927
$\alpha=0.45$ 2.65416 3.50244 1.62916 2.23837 1.38455 1.86748 1.27639 1.68194 1.21559 1.46514
$\alpha=0.5$ 2.72474 3.53269 1.65068 2.26026 1.39672 1.88259 1.28479 1.6932 1.22198 1.51352
$\alpha=0.674$ 2.96783 3.57489 1.72274 2.33008 1.43708 1.93236 1.31253 1.73039 1.24305 1.5618
(a) for $\alpha=0.674$
(b) for $\alpha=0.5$
(c) for $\alpha=0.45$
(d) for $\alpha=0.2$
Figure 3: The geometrical shape of the shadow in celestial plane of the higher-dimensional MOG dark compact object with $M=1$.
Plotting the shadow circles of the MOG dark compact object with extra dimensions for various values of $\alpha$ and $D$ is possible using the information given in Table 1. Figure 3 shows the geometrical shapes of the MOG dark compact object’s shadow with extra dimensions in celestial coordinates for various values of $\alpha$ and $D$. As one can see it shows plots for fixed values of $\alpha$ and illustrates how the black hole’s shadow shapes get smaller as dimension increases for a certain value of $\alpha$. Therefore, by reducing the size of shadow’s geometrical shape, the extra dimensions have a major impact on the black hole’s shadow.
Additionally, the MOG dark compact object’s shadow circles with extra dimensions in the celestial coordinates for various values of $\alpha$ are visible in Fig.4. As one can see for $D=4$ in Fig. 4a and $D=5,6$ in Figs 4b and 4c. As we can see in Fig. 4a for $D=4$, the size of the shadow circles get larger as the STVG parameter value increases. However, in Figs. 4b and 4c, the MOG dark compact object’s shadow size for $D=5,6$ with higher dimensions for each value of $\alpha$, approaching each other even as their sizes decrease relative to comparable ones for $D=4$. Additionally, the behavior for $D>6$ is obvious. Consequently, Fig. 4 demonstrates that the shadow circles corresponding to various values of $\alpha$ coincide on one another for $D\geq 6$. Consequently, with STVG theory, the effect of $\alpha$ on the higher-dimensional dark compact object’s shadow is amplifier.
(a) for $D=4$
(b) for $D=5$
(c) for $D=6$
Figure 4: The geometrical shape in celestial plane of the shadow of MOG dark compact object with higher dimensions with $M=1$.
III.3 Energy Emission rate
Hawking radiation is a phenomenon that allows black holes to radiate. The absorption cross-section typically oscillates around a limiting constant $\sigma_{lim}$ at very high energies. However, the absorption cross-section advances in the direction of the black hole shadow for an observer at a very far distance Wei:2013kza ; Belhaj:2020rdb . It is possible to demonstrate that $\sigma_{lim}$ is roughly equivalent to the area of the photon sphere, which may be expressed as follows in arbitrary dimension Wei:2013kza ; Li:2020drn ; Decanini:2011xw
$\sigma_{lim}\approx\frac{\pi^{\frac{n-2}{2}}}{\Gamma\left[\frac{n}{2}\right]}R_{s}^{n-2}\,.$ (40)
Thus, the complete form of the energy emission rate of higher-dimensional black holes can be read as
$\frac{d^{2}E(\varpi)}{d\varpi dt}=\frac{2\pi^{2}\sigma_{lim}}{e^{\frac{\varpi}{T}}-1}\varpi^{D-1}\,,$ (41)
where $\varpi$ is the emission frequency and $T$ is the Hawking temperature of the black hole.
In order to obtain the values of the Hawking temperature of the MOG dark compact object with extra dimensions for various values of $\alpha$ and $D$, we now use the values of $r_{eh}$ from Table 1. Next, by inserting the values of $\sqrt{\eta+\xi^{2}}$ into Eq. (40), one can determine the values of $\sigma_{lim}$ for various values of $\alpha$ and $D$ which correspond to the higher-dimensional MOG dark compact object. In order to obtain the energy emission rate in terms of various values of $\alpha$ and $D$ associated with the higher-dimensional MOG dark compact object, one can finally use the values of the Hawking temperature and $\sigma_{lim}$ in Eq. (41).
(a) for $\alpha=0.45$
(b) for $\alpha=0.2$
(c) for $D=4$
(d) for $D=5$
Figure 5: The energy emission rate as a function of $\varpi$ for different values of $D$ and $\alpha$.
Fig. 5 shows the energy emission rate as a function of the emission frequency for the higher-dimensional MOG dark compact object. The energy emission rate of MOG dark compact object with extra dimensions greatly increases by growing $D$, as shown in Figs. 5a and 5b, at a fixed value of $\alpha$. Consequently, we discover that in STVG with higher dimensions, extra dimensions speed up the MOG compact object’s evaporation. Also, the impact of the STVG parameter is to amplify the energy emission rate of MOG dark compact object, which causes it’s evaporation to accelerate.
III.4 Deflection Angle
We aim to give a framework for examining the deflection angle of higher-dimensional black holes in the line element (8). We intend to apply the Gauss-Bonnet theorem Gibbons:2008rj ; Arakida:2017hrm . To begin, we compute the optical metric on the equatorial hyperplane $\theta_{i}=\pi/2$ using the line element (8). On this hyperplane, we set $d\theta_{D-2}^{2}=d\phi^{2}$ to find
$ds^{2}=-h(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\phi^{2}\,.$ (42)
Then, for the considered null geodesics for which $ds^{2}=0$, the optical metric reads as follows
$dt^{2}=\frac{dr^{2}}{h(r)f(r)}+\frac{r^{2}}{h(r)}d\phi^{2}\,.$ (43)
For this optical metric, we can calculate the Gaussian optical curvature $K=\frac{\bar{R}}{2}$ in which $\bar{R}$ is the Ricci scalar of the metric (43) as follows
$K=\frac{2\,rh(r)f(r)h^{\prime\prime}(r)-2\,rf(r)h^{\prime 2}(r)+h(r)h^{\prime}(r)\left\{rf^{\prime}(r)+2f(r)\right\}-2f^{\prime}(r)h^{2}(r)}{2\,rh(r)}\,.$ (44)
In order to calculate the deflection angle, one should consider a non-singular manifold $\mathcal{D}_{\tilde{R}}$ with a geometrical size $\tilde{R}$ to employ the Gauss-Bonnet theorem, so that Gibbons:2008rj ; Arakida:2017hrm
[…]
where $dS=\sqrt{\bar{g}}drd\phi$ and $dt$ are the surface and line element of the optical metric (43), respectively, $\bar{g}$ is the determinant of the optical metric, $k$ denotes the geodesic curvature of $\mathcal{D}_{\tilde{R}}$, and $\varphi_{i}$ is the jump (exterior) angle at the $i$-th vertex, and also, $\zeta(\mathcal{D}_{\tilde{R}})$ is the Euler characteristic number of $\mathcal{D}_{\tilde{R}}$. One can set $\zeta(\mathcal{D}_{\tilde{R}})=1$. Then, considering a smooth curve $y$, which has the tangent vector $\dot{y}$ and acceleration vector $\ddot{y}$, the geodesic curvature $k$ of $y$ can be defined as follows where the unit speed condition $\tilde{g}\left(\dot{y},\dot{y}\right)=1$ is employed
$k=\tilde{g}\left(\nabla_{\dot{y}\dot{y},\ddot{y}}\right)\,,$ (46)
[…]
Finally, using the straight light ray approximation $r(\phi)=\xi/\sin\phi$, the Gauss-Bonnet theorem results in the following expression to calculate the deflection angle (for more details, see Refs. Gibbons:2008rj ; Arakida:2017hrm and references therein)
[…]
Applying Eq. (9) into Eqs. (43) and (44), yields the Gaussian optical curvature for the higher dimensional MOG dark compact object
$K\approx\frac{21M(D-3)}{2}\bigg{\{}\Big{[}\frac{2((D-2)\alpha+\frac{D}{3}-\frac{1}{3})(1+\alpha)Mr^{(5-2D)}}{7}\Big{]}-\Big{[}\frac{2(1+\alpha)(D-2)r^{(2-D)}}{21}\Big{]}+\Big{[}\alpha M^{2}r^{(8-3D)}(D-\frac{20}{7})\Big{]}\bigg{\}}$ (49)
Furthermore, the surface element of the optical metric (42) for the higher-dimensional MOG dark compact object in correspondance with the metric coefficients (9) can be found approximately as follows
$dS=\sqrt{\bar{g}}\,drd\phi=\frac{r}{h(r)\sqrt{f(r)}}dtd\phi\approx rdrd\phi\,.$ (50)
Now, employing Eqs. (49) and (50) in the deflection angle formula (48) leads to the deflection angle of the higher-dimensional MOG dark compact object as follows
[…]
(a) for $\alpha=0.45$
(b) for $D=5$
(c) for $D=6$
Figure 6: The behavior of the deflection angle of the higher-dimensional MOG dark compact object in terms of $\xi$ for different values of $D$ and $\alpha$..
The behavior of the deflection angle of the higher-dimensional MOG dark compact object is illustrated in Fig. 6 for different values of $D$ with respect to $\alpha=0.45$ in Fig. 6a and for different values of $\alpha$ with respect to $D=4,5$ in Figs. 6b and 6c, respectively. Fig. 7a, shows us that decreasing the value of the impact parameter $\xi$ results in extremely increasing the deflection angle of the MOG dark compact object. From Fig. 6a, we see that for a fixed value of STVG parameter, the deflection angle of the MOG dark compact object reduces by growing the number of extra dimensions. In Fig. 6b and Fig. 6c we see that for a fixed value of $D$, the deflection angle of the black hole decreases by increasing STVG parameter $\alpha$.
IV Constraints from EHT observations of M87*
Here we aim to compare the deduced shadow radius of the higher-dimensional MOG dark compact object with the shadow size of supermassive black hole, M87* captured by EHT. Within $1$-$\sigma$ (68%) confidence level, one can find that the shadow size of M87* supermassive black hole captured by EHT lies within the interval EventHorizonTelescope:2021dqv
$4.31\leq R_{s,M87^{*}}\leq 6.08\,.$ (52)
Comparing this with the shadow size of the higher-dimensional MOG dark compact object enables us to constrain the STVG parameter $\alpha$.
Figure 7: The shadow radius of the extra-dimensional MOG dark compact object, compared to the shadow size of $M87^{*}$ captured by EHT within $1-\sigma$ confidence level versus the parameter $\alpha$. The colored area is excluded region, which is inconsistent with the EHT data, while the white region is the $1-\sigma$ confidence level of EHT data.
Figure 7 indicates the behavior of the shadow radius of the higher-dimensional MOG dark compact object in comparison with the EHT’s shadow size of M87* within $1$-$\sigma$ uncertainties given in Eq. (52) versus the STVG parameter $\alpha$. In this figure, the white (unshaded) region denotes the $1$-$\sigma$ confidence level while the brown (shaded) areas are the excluded regions, which are incompatible with the EHT observations associated with shadow radius of M87*. From this figure, we see that the shadows of the higher-dimensional MOG dark compact object associated with $D=5,6,7$ are incompatible with the observations of EHT. However, the shadow of the MOG dark compact object with $D=4$ lies within the $1$-$\sigma$ confidence level of EHT data, so that in the range $0.2790\leq\alpha<$ 0.279 the shadow radius of the four dimensional MOG dark compact object has a good consistency with EHT observations. Moreover, like Table 1 and Fig. 3a, we see from Fig. 7 that increasing the STVG parameter value leads to amplify the shadow radius of the four dimensional MOG dark compact object. For $D\geq 5$ the impact of the STVG parameter in comparison with the extra dimension effect is somehow negligible and the effect of extra dimensions is dominated as it is obvious in Fig. 7
V Acceleration bounds for higher-dimensional MOG dark compact object
As an important issue in this setup, in this section we study the correlation between the shadow radius of the higher-dimensional MOG dark compact object and the acceleration bounds on radial linear uniformly accelerated (LUA) trajectories, i.e. the curved spacetime equivalent of the hyperbolic Rindler trajectories in a flat spacetime. The Rindler trajectories in flat spacetime can be described in a covariant way using the Letaw-Frenet equations, which involve three geometric scalars for a curve. For these trajectories, only the curvature scalar is non-zero and constant, and it is equal to the acceleration magnitude. The other two scalars, torsion and hyper-torsion, are zero. This covariant definition can be extended to curved spacetime, and the resulting curves with constant curvature scalar and zero torsion and hyper-torsion scalars are locally hyperbolic and linear in a local inertial frame at any point on the curve. From the linearity condition for LUA, the uniform acceleration can be expressed as the following equation
$u^{j}\nabla_{j}a^{i}-a^{2}u^{i}=0,$ (53)
where $a^{i}$ and $u^{i}$ are respectively the acceleration and velocity vectors along the trajectory. By solving the above equations for a given spacetime, one can find the set of LUA trajectories with the constant magnitude of acceleration $|a|$. Therefore, we obtain LUA trajectory consistent with the constraint of the equations (53) by considering the higher-dimensional MOG dark compact object metric (9)
$\frac{dt}{d\lambda}=f^{-1}(r)(|a|r+h),$ (54)
$\frac{dr}{d\lambda}=\pm\sqrt{(|a|r+h)^{2}-f(r)},$ (55)
where $\lambda$ is the proper time along the trajectory, $|a|$ is the uniform acceleration magnitude and $h$ is defined as a boundary data. In the case where $h=0$, the radial LUA trajectories follow the relation
$a^{2}r^{2}-1-\frac{M^{2}r^{2(D-3)}\alpha(1+\alpha)}{(r^{2(D-3)}+M^{2}\alpha(1+\alpha))^{2}}+\frac{2Mr^{2(D-3)}(1+\alpha)}{(r^{2(D-3)}+M^{2}\alpha(1+\alpha))^{\frac{3}{2}}}=-V_{eff}$ (56)
in which $|a|\neq 0$. The above equation specifies the kinematics of the radial LUA trajectory. As a condition at turning point $r_{t}$ of this trajectory, the radial velocity should be zero which constrains the effective potential to be zero. Hence,
$-V_{eff}=a^{2}r_{t}^{2}\left(r_{t}^{2D-6}+\alpha(\alpha+1)M^{2}\right)^{3/2}-\left(r_{t}^{2D-6}+\alpha(\alpha+1)M^{2}\right)^{3/2}-\frac{\alpha(\alpha+1)M^{2}r_{t}^{2D-6}}{\sqrt{a^{2D-6}+\alpha(\alpha+1)M^{2}}}+2(\alpha+1)Mr_{t}^{2D-6}=0$ (57)
At the bound $|a|=|a|_{b}$, the turning point is an extremum of the above equation, thus we have
$\begin{split}0=\frac{1}{r_{b}^{7}}&\bigg{\{}r_{b}^{2D}\Big{[}(3D-7)a_{b}^{2}r_{b}^{2}-3D+9\Big{]}\sqrt{r_{b}^{2D-6}+\alpha(\alpha+1)M^{2}}+4(\alpha+1)(D-3)Mr_{b}^{2D}+\frac{1}{\left(r_{b}^{2D}+\alpha(\alpha+1)M^{2}r_{b}^{6}\right)^{2}}\\ &\times\bigg{(}\alpha(\alpha+1)M^{2}r_{b}^{6}\sqrt{r_{b}^{2D-6}+\alpha(\alpha+1)M^{2}}\Big{[}2\alpha(\alpha+1)M^{2}r_{b}^{2D+6}\left(2a_{b}^{2}r_{b}^{2}-D+3\right)+r_{b}^{4D}\left(2a_{b}^{2}r_{b}^{2}-D+3\right)\\ &+2\alpha^{2}(\alpha+1)^{2}M^{4}a_{b}^{2}r_{b}^{14}\Big{]}\bigg{)}\bigg{\}}\end{split}$ (58)
By solving equations (58) and (57) simultaneously, one can find the following equations for the turning point
$|a|_{b}=\frac{\sqrt{\frac{r_{b}^{4D}+M^{4}r_{b}^{12}\alpha^{2}(1+\alpha)^{2}+Mr_{b}^{(6+2D)}(1+\alpha)(3M\alpha-2\sqrt{r_{b}^{(2D-6)}+M^{2}\alpha(1+\alpha)})}{r_{b}^{12}\sqrt{r_{b}^{(2D-6)}+M^{2}\alpha(1+\alpha)}}}}{r_{b}(r_{b}^{(2D-6)}+M^{2}\alpha(1+\alpha))^{\frac{3}{4}}}$ (59)
$\begin{split}0=-\frac{2}{\left(r^{2D}+\alpha(\alpha+1)M^{2}r^{6}\right)^{3}}&\bigg{\{}(\alpha+1)Mr^{4D+6}\left(\alpha(D+1)M-(D-1)\sqrt{r^{2D-6}+\alpha(\alpha+1)M^{2}}\right)\\ &+\alpha(\alpha+1)^{2}M^{3}r^{2(D+6)}\left(2(D-4)\sqrt{r^{2D-6}+\alpha(\alpha+1)M^{2}}-\alpha(D-7)M\right)\\ &+r^{6D}+\alpha^{3}(\alpha+1)^{3}M^{6}r^{18}\bigg{\}}\end{split}$ (60)
Since the equations (60) and (34) are the same, the bound on the distance of closest approach $r_{b}$ will be equal to the radius of photon sphere $r_{0}$. By setting $r_{b}=r_{0}$ and doing some calculations, one can find out that the equation (59) is related to the shadow radius of the higher-dimensional MOG dark compact object as follows
$|a|_{b}=\frac{1}{R_{s}}$ (61)
In what follows we provide numerical results of the acceleration bound for LUA trajectory in a higher-dimensional MOG dark compact object background. As we expected, the acceleration bound $|a|_{b}$ is the reverse of the value of the shadow radius given in subsection III.2.
Table 2: The numerical values of acceleration bound $|a|_{b}$ for LUA trajectory in a higher-dimensional MOG dark compact object spacetime for some different values of the parameters $\alpha$ and $D$.
$\alpha$ $|a|_{b}$
$D=4$ $D=7$ $D=8$
$0.674$ $0.139761$ $0.379497$ $0.449613$
$0.45$ $0.151589$ $0.395799$ $0.446273$
$0.2$ $0.170696$ $0.418589$ $0.50117$
VI Concluding Remarks
In this study, according to string theory, braneworld models, and AdS/CFT correspondence, we motivated to obtain a higher-dimensional MOG dark compact object solution. We investigated the behaviors of the corresponding MOG dark compact object’s shadow and deflection angle. Our main goal was to discover how extra dimensions and the STVG parameter affect the shadow of the black holes. To do this, we first developed a higher-dimensional metric for a dark compact object in STVG theory, and then provided the required general formalism to study the shadow behavior of these higher-dimensional black holes utilizing the Hamilton-Jacobi approach and Carter method to formulate the null geodesics around them and derive the corresponding effective potentials. Next, we introduced the celestial coordinates to specify the shadow shape of the higher-dimensional STVG black holes on the observer’s sky. We also derived the energy emission rate and deflection angle formulas in this higher-dimensional STVG scenario. Then, employing the constructed framework, we studied the shadow behavior, deflection angle, and energy emission rate of MOG dark compact object with extra dimensions. We computed and analyzed the significant impacts of the STVG parameter and extra dimensions on the shadow, deflection angle, and energy emission rate of the black holes within this setup. Moreover, we constrain these parameters by comparing the shadow size of M87* from EHT observations with the shadow radius of the higher-dimensional MOG dark compact object. Finally, the issue of acceleration bounds in higher-dimensional MOG dark compact object has been studied in this context in confrontation with EHT data of M87*.
For the higher-dimensional MOG dark compact object, we discovered that for a fixed value of the STVG parameter $\alpha$, the shadow size of the MOG dark compact object decreases with increasing the number of extra dimensions $D$. Also, when the STVG parameter value increases with a fixed $D$, the shadow size of the MOG dark compact object increases. Also, we saw that for a fixed value of $\alpha$, the energy emission rate of the MOG dark compact object with extra dimensions extremely increases by growing $D$. Also, growing the STVG parameter increases the energy emission rate, not as mush as the extra dimensions. Therefore, we found that extra dimensions accelerate the evaporation of the MOG dark compact object with higher dimensions. Then, using the Gauss-Bonnet theorem, we have calculated the leading terms of the deflection angle in the weak-limit approximation. We have discussed the impact of the STVG parameter and the extra dimensions on this optical quantity. We noticed that for a fixed value of the STVG parameter, the deflection angle of the MOG dark compact object with extra dimensions reduces by growing the number of dimensions. Also, for a fixed value of $D$, the deflection angle of the MOG dark compact object decreases by increasing the STVG parameter $\alpha$. Furthermore, by comparing the shadow radius of the MOG dark compact object with M87* shadow released by EHT, we observed that only the shadow of four dimensional MOG dark compact object with $0.2790\leq\alpha<$ 0.279 lies in the $1$-$\sigma$ uncertainties of EHT data.
In summary, we can came to conclusion that the shadow of higher-dimensional MOG dark compact object is characterized by the extra dimensions in addition to the STVG parameter. In this regard, the extra dimensions within STVG theory affect the shadows of the dark compact object by reducing their size, significantly. On the other hand, we concluded that it seems possible to detect the effects of the STVG theory via EHT. The key point here is that from Fig. 7, one can expect that the limit $\alpha<0.279$ can be apparently observed from the shadow of black holes captured by EHT. These outcomes may lead to the possibility of testing the modified theories of gravity, and their solutions and their parameters by employing astrophysical observations.
Acknowledgements.
The authors would like to thank John W. Moffat for their fruitful comments and discussions, which improved the quality of the paper, considerably.